Hello I need help on finding the circumference of a circle whose equation is: x^2 + y^2 = 16, the symbol ^ means squared.
Answer by stanbon(75887)

Hello I need help on finding the circumference of a circle whose equation is: x^2 + y^2 = 16, the symbol ^ means squared.
------
radius = sqrt(16) = 4
-----
Circumference = 2*pi*r = 8*pi is approximately 25.13
=========================
Cheers,
Stan H.